Find two numbers such that their sum is 10 and their product is 22.
Answer #1
A)x+y=10 B)xy=22 . y=22/x from B) . x+(22/x)=10 substitute 22/x for y in A) x^2+22=10x multiply x times each side. x^2-10+22=0 Quadratic formula: x=5+sqrt(3) or x=5-sqrt(3) . When x=5+sqrt(3) y=10-(5+sqrt(3))=10-5-sqrt(3)=5-sqrt(3) When x=5-sqrt(3) y=10-(5-sqrt(3))=10-5+sqrt(3)=5+sqrt(3) . Check: 5+sqrt(3)+5-sqrt(3)=10 (5+sqrt(3))*(5-sqrt(3))=25-3=22 Ed
answered by: kiren